# The frustrated Ising model on the honeycomb lattice: Metastability and universality

### v1

@@ III.5 Simulation details
Rejection-free PA – We implemented rejection-free PA by means of parallel CPU code using the OpenMP interface running on single compute nodes. Our implementation of the PA framework is based on the open source code provided by the authors of Ref. [45]. As the $n$-fold way update algorithm is inherently difficult to parallelize, we employ one thread per replica. We use R=20  [^0] R=20\,000 as population size, and an adaptive cooling protocol with a target histogram overlap $\alpha^{*}=0.8$. When the number of sweeps is chosen adaptively, we target an effective population size $R_{\text{eff}}(E)/R_{i}=\gamma^{*}=0.9$. Otherwise, $n_{\text{f}}$ is chosen constant throughout one simulation and as large as necessary or as large as feasible (whichever is lower). We use the nearest-integer resampling method [40] in step 22(b) of the PA algorithm.
Metropolis PA – As a baseline comparison for the proposed adaptive $n$-fold way PA algorithm we employ a graphics processing unit (GPU) implementation of the PA framework based on the code provided in Ref. [31]. In addition to the parallelism on the replica-level also present in rejection-free PA, the Metropolis algorithm is parallelized by means of a domain decomposition into 4 sublattices (see Appendix B). This highly parallel setting is particularly well-suited for the execution on GPUs. Also here, we use a population size of R=20  [^0] R=20\,000 and the same adaptive temperature protocol. Again, we use the nearest-integer resampling method in step 22(b).
@@ IV.1 Overview
Figure 2: Phase diagram in the $J_{2}-T$ plane. The square symbols correspond to the pseudocritical temperature obtained from the peak locations of the specific heat for $L=48$ systems. Triangles denote the exactly known values for $J_{2}=0$ and $J_{2}=-1/4$. The solid colored lines show effective-field theory (EFT) approximations for various cluster sizes $n$ from Ref. [18], with the endpoints indicating the predicted tricritical locations. The black solid line shows exact pseudocritical points for an $L=4$ system obtained by exact enumeration. Blue circles show previous MC results from Ref. [20]. Green diamonds denote FSS extrapolations for $T_{c}$ from our previous work in Ref. [24] using MC data. The dashed line is merely a guide to the eye. Note that the EFT lines for $n>1$ do not exist below a certain value of $J_{2}$, where they (erroneously) predict a tricritical point.
@@
Both our estimates (squares) and those from Ref. [20] (circles) are obtained through the locations of the specific-heat peaks for $L=48$. For $J_{2}=-0.1$, 0.2-0.2, 0.21-0.21, and 0.22-0.22 we also show the estimates for $T_{c}$ of our recent work based on the analysis of partition function zeros as reported in Ref. [24], which are only slightly lower than these pseudocritical temperatures. Our results of the present work (obtained from rejection-free PA simulations with adaptive sweep schedules) are in very good agreement with previous data. The simulations of Ref. [20] failed to equilibrate for $J_{2}<-0.2$ due to very long relaxation times. In Ref. [24], we employed PA with the same adaptive sweep protocol as in the present study, however, using the Metropolis update. This approach was already for $J_{2}=-0.22$ quite costly, and failed for even lower values. Here, we can equilibrate $J_{2}\geq-0.23$ thanks to the rejection-free Monte Carlo updates [32] (see Sec. III.3). By means of FSS (see Sec. IV.3) we show that also for $0.23\leq J_{2}<-0.2$ the critical exponents are consistent with the Ising ones.
@@
One reason for the slow relaxation of the Metropolis dynamics for lower values of $J_{2}$ is that the transition takes place at a low temperature: assuming typical $\Delta E$’s to be of the same order as for other $J_{2}$, this results in small acceptance rates, and therefore unavoidably in slower dynamics. This effect is reduced by using the rejection-free update (see Sec. III.3). A slightly more subtle effect also contributing to this slow relaxation is that paramagnetic structures formed at the same value of $\beta/\beta_{c}$ are more stable for lower values of $J_{2}$; see Fig. 3 showing configuration snapshots for $J_{2}=0$,0.1-0.1,0.2-0.2, and 0.23-0.23 for $\beta=0.6\beta_{c}$. While the configurations in (a) and (b) are very noisy, (c) and (d) show a clear domain picture with smooth boundaries. Using single spin-flip dynamics, overturning such domains is only possible by passing through configurations of much higher energy, rendering these states rather stable (albeit only metastable below the critical temperature). The reason why also the rejection-free approach fails to overturn these domains for low enough values of $J_{2}$, is that it “flickers” between various excited states, that is, one effectively stays in the same subset of spin configurations for a long time. This is illustrated using the example of a single overturned hexagon later in Sec. IV.4.
@@ IV.2 Effect of the update and sweep schedule on equilibration
Figure 4: Results from PA simulations using Metropolis updates with $\theta=2000$ (resp. $\theta=500$ for $J_{2}=0$) and $L=48$. Top panel: Magnetization per spin $m$ for various coupling strengths. For $J_{2}\gtrsim-0.2$ we obtain good agreement with the rejection-free data (not shown here). For $J_{2}<-0.22$ the simulations fail to reach the fully ordered state at low temperatures. Bottom panel: Replica-averaged family size $\rho_{t}$. As expected [30] $\rho_{t}$ is monotonously increasing during the annealing process with decreasing temperature. The failure to equilibrate for $J_{2}<-0.2$ is reflected in a sharp increase in $\rho_{t}$ near the corresponding critical temperatures $T_{c}$ for $J_{2}\leq-0.2$.
@@
Figure 5: Overview over the thermodynamic quantities measured as functions of temperature for $J_{2}\in\{-0.2,-0.21,-0.22,-0.23\}$ for a system of linear size $L=48$ from PA simulations using $n$-fold way updates. Top left: Energy per spin (dashed lines correspond to the ground-state energy), Top right: Magnetization per spin. Bottom left: Specific heat. Bottom right: Magnetic susceptibility.
@@
Figure 6: Demonstration of efficacy of rejection-free PA with an adaptive number of sweeps at each temperature using $L=48$. Top panel: Inverse of the Metropolis acceptance probability (which is proportional to the speedup achieved by the $n$-fold way update) as a function of inverse temperature. $1/p_{\text{acc}}$ grows rapidly in $\beta$ throughout with a marked increase around criticality (colored, shaded regions). Bottom panel: Equivalent number of Metropolis attempts and number of flips made at each inverse temperature using the adaptive sweep protocol for $J_{2}=-0.23$. For $\beta>5$ one notes a significant increase in the number of flips. In this range, the inverse temperatures used in the PA run are indicated as $x$-ticks at the top of the plot.
The reason why the rejection-free approach improves equilibration is readily understood by considering the Metropolis acceptance rates $p_{\text{acc}}$ as a function of inverse temperature $\beta$, which across the critical regime decrease at least exponentially in $\beta$ before reaching the asymptotic behavior $p_{\text{acc}}=\exp(-(6+12J_{2})\beta)$ for large $\beta$ (top panel in Fig. 6). The colored, shaded areas indicate the critical regimes for the different $J_{2}$. Whereas for $J_{2}=-0.21$ still 1 in 100 spin flip attempts is accepted, it is only 1 in 5000 for $J_{2}=-0.23$, rendering the 2000 sweeps per temperature from the Metropolis PA simulations discussed above much too few to decorrelate spin configurations between temperature steps. In fact, $1/p_{\text{acc}}$ is directly proportional to the computational speed-up achieved by using the rejection-free approach as opposed to normal Metropolis. Note that the shown acceptance rates were obtained already using the adaptive sweep schedule (see Sec. III.4). (Since the acceptance probability is a thermodynamic observable, it is agnostic to the choice of the sweep schedule.)
@@
The bottom panel of Fig. 6 shows for $J_{2}=-0.23$ the number of $n$-fold way flips carried out per site and temperature, as well as the equivalent number of spin-flip attempts in the Metropolis scheme using the same simulation data as in the top panel. Specifically, the adaptive sweep criterion of $R_{\text{eff}}(E)>0.9R_{i}$ was imposed. As one can see, for small $\beta$ only very few flips are required to decorrelate the replicas (25 was chosen as a lower bound), whereas for $\beta\approx 5$ and higher many more flips are required. The adaptive sweep criterion has two benefits: On the one hand, the simulation quickly passes through the easily-to-equilibrate high temperatures, and on the other hand, most time is spent in the much harder-to-equilibrate critical regime and at lower temperatures. Note also that the adaptively chosen temperatures (using $\alpha^{*}=0.8$) are denser in the critical region (indicated by the $x$-ticks at the top of the plot for $\beta>5$). Within this region, in total $5\times 10^{8}$ Metropolis MCS are carried out per replica, which corresponds to only $10^{5}$ rejection-free spin flips per site and replica (making up 55% of the overall number of spin flips of this PA run). In principle, one can use adaptive sweeps also with regular Metropolis, which in fact is more efficient than without an adaptive schedule. However, with such low acceptance rates it takes a very long time to equilibrate at low temperatures (see equivalent Metropolis attempts in Fig. 6). Our previous simulations of Ref. [24] were carried out using regular Metropolis PA combined with the adaptive sweep protocol. In this study equilibrating systems with $J_{2}=-0.22$ was difficult but still possible (also note the smaller system sizes there). For $J_{2}=-0.23$, on the other hand, the adaptive sweep protocol using the Metropolis update fails to reach temperatures below the critical one within a reasonable time even for small systems. Note that the visible jumps in the lower panel are artifacts of the discretization used in the sweep protocol.
Lastly, the replica-averaged family size $\rho_{t}$ (Fig. 7) is studied for two rejection-free runs, one with adaptive sweep schedule and one without. First, the adaptive one was run, and then the average number of flips per temperature was calculated, to ascertain that both runs use approximately the same CPU time. For small $\beta$ both curves have similar values, indicating that at high temperature indeed very few spin flips are required for equilibration. Close to criticality $\rho_{t}$ of the run without an adaptive sweep schedule then increases rapidly, whereas it continues without a steep increase otherwise. Thus, at the same computational cost much better equilibration is to be expected from the adaptive run.
@@ IV.3 Finite-size scaling analysis
Figure 8: Finite-size scaling fits using Eq. (4b) and Eqs. (4d)-(4f) for $J_{2}=-0.21$ and 0.23-0.23 for system sizes up to $L=128$. Solid lines correspond to the fit ranges used to obtain the exponents displayed in Table 1. Dotted lines are extrapolations of the fits to values of $L$ outside the fitting range.
@@
We now turn to studying the critical behavior of the ferromagnetic ordering by means of a finite-size scaling (FSS) analysis [34] for $J_{2}=-0.21$, 0.22-0.22, and 0.23-0.23. All simulations for $L=12$, $16$, $24$, $32$, $48$, $64$, $88$, and $128$ were carried out using rejection-free PA with the adaptive sweep schedule.
@@
Figure 8 shows for $J_{2}=-0.21$ and $J_{2}=-0.23$ the peak heights of $\chi$, $\mathrm{d}\hskip-0.50003pt\ln|m|$, and $\mathrm{d}\hskip 0.50003pt|m|$, as well as $m$ evaluated at the inverse temperature of the $\mathrm{d}\hskip-0.50003pt\ln|m|$ peak; the values were determined through (single) histogram reweighting [51, *Ferrenberg1989] of the PA data  [^2] Note that the simulation temperature closest to the corresponding peak location is used as reference temperature in histogram reweighting. Alternatively, it would also be possible to use the populations from all temperatures in a multi-histogram approach, see Ref. [31].. The finite-size scaling relations (4a)-(4g) only hold asymptotically, and in principle one expects (unknown) correction terms that vanish as $L\rightarrow\infty$. To take these into account, system sizes less than $L=32$ were discarded in the fits, such that all fits have a quality of fit $Q>0.1$. While we will elaborate on the fitting results in more detail below in the discussion of Table 1, note that the obtained values for $1/\nu$ of 0.96(1) for $J_{2}=-0.21$, and 0.98(1) for $J_{2}=-0.23$ are consistent with the expected Onsager value of 1, albeit not fully within error bars.
@@
Table 1: Critical exponents and corresponding estimates for the inverse critical temperature $\beta_{c}$ determined through FSS fits for $J_{2}=-0.21$, $J_{2}=-0.22$, and $J_{2}=-0.23$. (Fits for $J_{2}=-0.21$, and $J_{2}=-0.23$ are shown in Figs. 8 and 9.)
@@
$\beta/\nu$ Eq. (4f) 0.117(13) 0.115(12) 0.115(13) $0.125$
Eq. (4g) 0.114(4) 0.119(4) 0.135(9)
denotes results from fits at $\beta_{c}$. Otherwise, fits are at pseudocritical points. For $J_{2}=-0.21$ and 0.22-0.22 our estimates for $\beta_{c}$ of Ref. [24] are used, and for $J_{2}=-0.23$ the estimate from the scaling of $\beta_{\max}^{(\mathrm{d}\hskip-0.40974pt\ln|m|)}$ is used.
denotes results from Ref. [24] obtained from the scaling of the partition function zeros. $\gamma/\nu$ is calculated from the fitting result of the Lee-Yang zeros, and $1/\nu$ from the Fisher zeros. In both cases the fitting range 248824-88 was selected such that the degrees of freedom are identical to the fits here.
@@
The obtained critical exponents and inverse critical temperatures are summarized in Table 1. Almost all obtained exponents are within one or two standard deviations of the expected Onsager exponents listed in the right column. This suggests that the model remains in the Ising universality class even rather close to the special point $J_{2}=-1/4$. One exception are the $\gamma/\nu$ exponents from the fit for the peaks of the magnetic susceptibility with subtraction. It is well known that using the magnetic susceptibility with subtraction (especially of the finite-size type $\langle|m|\rangle$) is subject to strong corrections (see, e.g., Refs. [54, 55, 56, 57]). The disconnected form $\chi^{\prime}$ for $\beta\leq\beta_{c}$ evaluated at $\beta=\beta_{c}$ commonly shows better scaling, as is also the case here. Nonetheless, even the $\gamma/\nu$ values from $\chi_{L}^{\prime}(\beta_{c}$) are about two to three standard deviations away from $1.75$. For $J_{2}=-0.21$ and 0.22-0.22 we use our $\beta_{c}$ estimates of Ref. [24], i.e., $\beta_{c}=3.2572(5)$ and $4.2702(6)$. For $J_{2}=-0.23$ we use $6.346(3)$ from the scaling of the location of $\mathrm{d}\hskip-0.50003pt\ln|m|$, i.e., $\beta_{\max}^{(\mathrm{d}\hskip-0.40974pt\ln|m|)}$. Similarly, evaluating the magnetization at $\beta_{\max}^{(\mathrm{d}\hskip-0.40974pt\ln|m|)}$ and at $\beta_{c}$ gives different estimates for the exponent ratio $\beta/\nu$. Table 1 also lists the results from the partition function zeros studied in Ref. [24] for $J_{2}=-0.21$ and 0.22-0.22, which give results compatible within error bars for $\gamma/\nu$, $1/\nu$, and $\beta_{c}$.

### v2

@@ III.5 Simulation details
Rejection-free PA – We implemented rejection-free PA by means of parallel CPU code using the OpenMP interface running on single compute nodes. Our implementation of the PA framework is based on the open source code provided by the authors of Ref. [45]. As the $n$-fold way update algorithm is inherently difficult to parallelize, we employ one thread per replica. We use R=20  [^0] R=20\,000 as population size, and an adaptive cooling protocol with a target histogram overlap $\alpha^{*}=0.8$. When the number of sweeps is chosen adaptively, we target an effective population size $R_{\text{eff}}(E)/R_{i}=\gamma^{*}=0.9$. Otherwise, $n_{\text{f}}$ is chosen constant throughout one simulation and as large as necessary or as large as feasible (whichever is lower). We use the nearest-integer resampling method [40] in step 22(b) of the PA algorithm. Error bars are obtained using Eq. (12).
Metropolis PA – As a baseline comparison for the proposed adaptive $n$-fold way PA algorithm we employ a graphics processing unit (GPU) implementation of the PA framework based on the code provided in Ref. [31]. In addition to the parallelism on the replica-level also present in rejection-free PA, the Metropolis algorithm is parallelized by means of a domain decomposition into 4 sublattices (see Appendix A). This highly parallel setting is particularly well-suited for the execution on GPUs. Also here, we use a population size of R=20  [^0] R=20\,000 and the same adaptive temperature protocol. Again, we use the nearest-integer resampling method in step 22(b).
@@ IV.1 Overview
Figure 2: Phase diagram in the $J_{2}-T$ plane. The square symbols correspond to the pseudocritical temperature obtained from the peak locations of the specific heat for $L=48$ systems. Triangles denote the exactly known values for $J_{2}=0$ and $J_{2}=-1/4$. The solid colored lines show EFT approximations for various cluster sizes $n$ from Ref. [18], with the end points indicating the predicted tricritical locations. The black solid line shows exact pseudocritical points for an $L=4$ system obtained by exact enumeration. Blue circles show previous MC results from Ref. [20]. Green diamonds denote FSS extrapolations for $T_{c}$ from our previous work in Ref. [24] using MC data. The dashed line is merely a guide to the eye. Note that the EFT lines for $n>1$ do not exist below a certain value of $J_{2}$, where they (erroneously) predict a tricritical point.
@@
Both our estimates (squares) and those from Ref. [20] (circles) are obtained through the locations of the specific-heat peaks for $L=48$. For $J_{2}=-0.1$, 0.2-0.2, 0.21-0.21, and 0.22-0.22 we also show the estimates for $T_{c}$ of our recent work based on the analysis of partition function zeros as reported in Ref. [24], which are only slightly lower than these pseudocritical temperatures. Our results of the present work (obtained from rejection-free PA simulations with adaptive sweep schedules) are in very good agreement with previous data. The simulations of Ref. [20] failed to equilibrate for $J_{2}<-0.2$ due to very long relaxation times. In Ref. [24], we employed PA with the same adaptive sweep protocol as in the present study, however, using the Metropolis update. This approach was already quite costly for $J_{2}=-0.22$ and failed for even lower values of $J_{2}$. Here we can equilibrate $J_{2}\geq-0.23$ thanks to the rejection-free Monte Carlo updates [32] (see Sec. III.3). By means of FSS (see Sec. IV.3) we show that also for $0.23\leq J_{2}<-0.2$ the critical exponents are consistent with the Ising ones.
@@
One reason for the slow relaxation of the Metropolis dynamics for lower values of $J_{2}$ is that the transition takes place at a low temperature: Assuming typical $\Delta E$’s to be of the same order as for larger $J_{2}$, this leads to small acceptance rates, and therefore unavoidably to slower dynamics. This effect is reduced by using the rejection-free update (see Sec. III.3). A slightly more subtle effect also contributing to this slow relaxation is that paramagnetic structures formed at the same value of $\beta/\beta_{c}$ are more stable for lower values of $J_{2}$; see Fig. 3 showing configuration snapshots for $J_{2}=0$,0.1-0.1,0.2-0.2, and 0.23-0.23 for $\beta=0.6\beta_{c}$. While the configurations in Figs. 3(a) and 3(b) are very noisy, Figs. 3(c) and 3(d) show a clear domain picture with smooth boundaries. Using single spin-flip dynamics, overturning such domains is only possible by passing through configurations of much higher energy, rendering these states rather stable (albeit only metastable below the critical temperature). The reason why also the rejection-free approach fails to overturn these domains for low enough values of $J_{2}$, is that it “flickers” between various excited states, that is, one effectively stays in the same subset of spin configurations for a long time. This is illustrated using the example of a single overturned hexagon later in Sec. IV.4.
@@ IV.2 Effect of the update and sweep schedule on equilibration
Figure 4: Results from PA simulations using Metropolis updates with $\theta=2000$ (respectively $\theta=500$ for $J_{2}=0$) and $L=48$. Top panel: Magnetization per spin $m$ for various coupling strengths. For $J_{2}\gtrsim-0.2$ we obtain good agreement with the rejection-free data (not shown here). For $J_{2}<-0.22$ the simulations fail to reach the fully ordered state at low temperatures. Bottom panel: Replica-averaged family size $\rho_{t}$. As expected [30] $\rho_{t}$ monotonously increases during the annealing process with decreasing temperature. The failure to equilibrate for $J_{2}<-0.2$ is reflected in a sharp increase in $\rho_{t}$ near the corresponding critical temperatures $T_{c}$ for $J_{2}\leq-0.2$.
@@
Figure 5: Overview over the thermodynamic quantities measured as functions of temperature for $J_{2}\in\{-0.2,-0.21,-0.22,-0.23\}$ for a system of linear size $L=48$ from PA simulations using $n$-fold way updates. The shaded areas indicate $2\sigma$ error environments of the data for $J_{2}\in\{-0.21,-0.22,-0.23\}$ (but note that in many cases they are hardly visible). Top left: Energy per spin (dashed lines correspond to the ground-state energy), Top right: Magnetization per spin. Bottom left: Specific heat. Bottom right: Magnetic susceptibility.
@@
Figure 6: Demonstration of efficacy of rejection-free PA with an adaptive number of sweeps at each temperature using $L=48$. Top panel: Inverse of the Metropolis acceptance probability (which is proportional to the speedup achieved by the $n$-fold way update) as a function of inverse temperature. $1/p_{\text{acc}}$ grows rapidly in $\beta$ throughout with a marked increase around criticality (colored, shaded regions). Bottom panel: Equivalent number of Metropolis attempts and number of flips made at each inverse temperature using the adaptive sweep protocol for $J_{2}=-0.23$. For $\beta>5$ one notes a significant increase in the number of flips. In this range, the inverse temperatures used in the PA run are indicated as $x$ ticks at the top of the plot.
The reason why the rejection-free approach improves equilibration is readily understood by considering the Metropolis acceptance rates $p_{\text{acc}}$ as a function of inverse temperature $\beta$, which across the critical regime decrease at least exponentially in $\beta$ before reaching the asymptotic behavior $p_{\text{acc}}=\exp(-(6+12J_{2})\beta)$ for large $\beta$ (top panel in Fig. 6). The colored, shaded areas indicate the critical regimes for the different $J_{2}$. While for $J_{2}=-0.21$ still 1 in 100 spin flip attempts is accepted, it is only 1 in 5000 for $J_{2}=-0.23$, rendering the 2000 sweeps per temperature from the Metropolis PA simulations discussed above much too few to decorrelate spin configurations between temperature steps. In fact, $1/p_{\text{acc}}$ is directly proportional to the computational speedup achieved by using the rejection-free approach as opposed to normal Metropolis. Note that the shown acceptance rates were obtained already using the adaptive sweep schedule (see Sec. III.4). (Since the acceptance probability is a thermodynamic observable, it is agnostic to the choice of the sweep schedule.)
@@
The bottom panel of Fig. 6 shows for $J_{2}=-0.23$ the number of $n$-fold way flips carried out per site and temperature, as well as the equivalent number of spin-flip attempts in the Metropolis scheme using the same simulation data as in the top panel. Specifically, the adaptive sweep criterion of $R_{\text{eff}}(E)>0.9R_{i}$ was imposed. As one can see, for small $\beta$ only very few flips are required to decorrelate the replicas (25 was chosen as a lower bound), whereas for $\beta\approx 5$ and higher many more flips are required. The adaptive sweep criterion has two benefits: On the one hand, the simulation quickly passes through the easily-to-equilibrate high temperatures, and on the other hand, most time is spent in the much harder-to-equilibrate critical regime and at lower temperatures. Note also that the adaptively chosen temperatures (using $\alpha^{*}=0.8$) are denser in the critical region (indicated by the $x$ ticks at the top of the plot for $\beta>5$). Within this region, in total $5\times 10^{8}$ Metropolis MCS are carried out per replica, which corresponds to only $10^{5}$ rejection-free spin flips per site and replica (making up 55% of the overall number of spin flips of this PA run). One can use adaptive sweeps also with regular Metropolis, which in fact is more efficient than without an adaptive schedule. However, with such low acceptance rates it takes a very long time to equilibrate at low temperatures (see equivalent Metropolis attempts in Fig. 6). As mentioned above, our previous simulations of Ref. [24] were carried out using this approach, and equilibrating systems with $J_{2}=-0.22$ was difficult but still possible (also note the smaller system sizes there). For $J_{2}=-0.23$, on the other hand, the adaptive sweep protocol using the Metropolis update fails to reach temperatures below the critical one within a reasonable time even for small systems. Note that the visible jumps in the lower panel are artifacts of the discretization used in the sweep protocol.
Last, the replica-averaged family size $\rho_{t}$ (Fig. 7) is studied for two rejection-free runs, one with adaptive sweep schedule and one without. First, the adaptive one was run, and then the average number of flips per temperature was calculated, to ascertain that both runs use approximately the same CPU time. For small $\beta$ both curves have similar values, indicating that at high temperature indeed very few spin flips are required for equilibration. Close to criticality, $\rho_{t}$ of the run without an adaptive sweep schedule then increases rapidly, whereas it continues without a steep increase otherwise. Thus, at the same computational cost much better equilibration is to be expected from the adaptive run.
@@ IV.3 Finite-size scaling analysis
We now turn to studying the critical behavior of the ferromagnetic ordering by means of an FSS analysis [34] for $J_{2}=-0.21$, 0.22-0.22, and 0.23-0.23. All simulations for $L=12$, $16$, $24$, $32$, $48$, $64$, $88$, and $128$ were carried out using rejection-free PA with the adaptive sweep schedule.
Figure 8: The Binder parameter $U_{4}$ in the vicinity of $T_{c}$ for different system sizes and for (a) $J_{2}=-0.21$, (b) 0.22-0.22, and (c) 0.23-0.23, respectively. Points denote the simulation temperatures and the connecting lines were calculated using single-histogram reweighting. The estimated critical temperatures $T_{c}(J_{2})$ are indicated by solid vertical lines and their margins of error by dashed vertical lines. The values for $1/T_{c}$ of $3.2572(5)$ and $4.2702(6)$ for $J_{2}=-0.21$ and 0.22-0.22, respectively, are taken from Ref. [24], while the value of $6.3455(14)$ for $J_{2}=-0.23$ is from Table 1. Similarly, the estimated values of $U^{*}=U_{4}(T_{c})$ (including their error margins) are shown by horizontal lines.
We begin the analysis by considering the Binder parameter $U_{4}$ and its crossings for different system sizes $L$ at $(T_{c},U^{*})$ for $J_{2}=-0.21$, 0.22-0.22, and 0.23-0.23 shown in Fig. 8. The data points with error bars correspond to the estimates at each annealing step, while the lines are extrapolations using (single) histogram reweighting [51, *Ferrenberg1989]. Since the individual simulation points are almost fully uncorrelated, the lines do not meet perfectly but instead small jumps of the order of the error bars are visible. As expected, the temperature curves of $U_{4}$ for different system sizes calculated in this way approximately cross at a single point $(T_{c},U^{*})$, providing a convenient and well-accessible estimate of the critical temperature $T_{c}$. Due to finite-size corrections, this crossing is not perfect and we have omitted the data for the two smallest system sizes considered, $L=12$ and $L=16$, as the crossings there were significantly shifted with respect to $T_{c}$. The vertical lines indicate the previously obtained critical temperature (with dashed lines as error margins), which for $J_{2}=-0.21$ and 0.22-0.22 are the values of Ref. [24], and for $J_{2}=-0.23$ is given by the value from Table 1 from the peak locations of $\mathrm{d}\hskip 0.50003pt|m|$ discussed below. In all three cases, the crossings are in good visual agreement with these values for $T_{c}$. For $U^{*}=\lim_{L\rightarrow\infty}U_{4}(T_{c},L)$ and for $J_{2}=-0.21$, 0.22-0.22, and 0.23-0.23, we obtain the values $0.6124(12)$, $0.6120(11)$, and $0.6117(11)$, respectively; see Appendix B for details.
Figure 9: Finite-size scaling fits using Eq. (5b) and Eqs. (5d)-(5g) for $J_{2}=-0.21$ and 0.23-0.23 for system sizes up to $L=128$. Solid lines correspond to the fit ranges used to obtain the exponents compiled in Table 1. Dotted lines are extrapolations of the fits to values of $L$ outside the fitting range.
@@
Figure 9 shows for $J_{2}=-0.21$ and $J_{2}=-0.23$ the peak heights of $\chi$, $\mathrm{d}\hskip-0.50003pt\ln|m|$, $\mathrm{d}\hskip 0.50003pt|m|$, and $\mathrm{d}U_{4}/\mathrm{d}\beta$, as well as $m$ evaluated at the inverse temperature of the $\mathrm{d}\hskip-0.50003pt\ln|m|$ peak; the values were determined through (single) histogram reweighting [51, *Ferrenberg1989] of the PA data  [^2] Note that the simulation temperature closest to the corresponding peak location is used as reference temperature in histogram reweighting. Alternatively, it would also be possible to use the populations from all temperatures in a multi-histogram approach, see Ref. [31].. The finite-size scaling relations (5a)-(5h) only hold asymptotically, and in principle one expects (unknown) correction terms that vanish as $L\rightarrow\infty$. To reduce the systematic error from omitting these finite-size corrections, system sizes $L<32$ were discarded in the fits, such that all fits have a quality of fit $Q>0.1$ – except for $\mathrm{d}U_{4}/\mathrm{d}\beta$ for which also $L=32$ had to be discarded in order to satisfy $Q>0.1$. While we will elaborate on the fitting results in more detail below in the discussion of Table 1, note that the obtained values for $1/\nu$ using Eq. (5d) of 0.96(1) for $J_{2}=-0.21$, and 0.98(1) for $J_{2}=-0.23$ are consistent with the expected Onsager value of 1, albeit not fully within error bars.
@@
Table 1: Critical exponents and corresponding estimates for the inverse critical temperature $\beta_{c}$ determined through FSS fits for $J_{2}=-0.21$, $J_{2}=-0.22$, and $J_{2}=-0.23$. (Fits for $J_{2}=-0.21$, and $J_{2}=-0.23$ are shown in Figs. 9 and 10.)
@@
$\beta/\nu$ Eq. (5g) 0.117(13) 0.115(12) 0.115(13) $0.125$
Eq. (5h) 0.114(4) 0.119(4) 0.135(9)
denotes results from fits at $\beta_{c}$. Otherwise, fits are at pseudocritical points. For $J_{2}=-0.21$ and 0.22-0.22 our estimates for $\beta_{c}$ of Ref. [24] are used, and for $J_{2}=-0.23$ the estimate from the scaling of $\beta_{\max}^{(\mathrm{d}\hskip-0.40974pt\ln|m|)}$ is employed.
denotes results from Ref. [24] obtained from the scaling of the partition function zeros. $\gamma/\nu$ is calculated from the fit result of the Lee-Yang zeros, and $1/\nu$ from the Fisher zeros. In both cases the fitting range 2424-$88$ was selected such that the degrees of freedom are identical to the fits here.
@@
The obtained critical exponents and inverse critical temperatures are summarized in Table 1. Almost all obtained exponents are within one or two standard deviations of the expected Onsager exponents listed in the right column. This suggests that the model remains in the Ising universality class even rather close to the special point $J_{2}=-1/4$. In contrast to the values of $1/\nu$ obtained from the FSS fits of the peak locations of $\mathrm{d}\hskip-0.50003pt\ln|m|$, the ones obtained via $\mathrm{d}U_{4}/\mathrm{d}\beta$ are particularly far from the expected value of 1. This may be due to $\mathrm{d}U_{4}/\mathrm{d}\beta$ exhibiting broader maxima than $\mathrm{d}\hskip-0.50003pt\ln|m|$ for the same $L$ in the present model (not shown), which also is reflected in much larger error bars for the obtained values of $\beta_{\max}^{\mathrm{d}U_{4}/\mathrm{d}\beta}$ in Fig. 10, particularly for smaller system sizes. Also note that the values for $\mathrm{d}U_{4}/\mathrm{d}\beta$ outside the fitting range in Fig. 9, i.e., data for $L<48$, fall well above the fitting line, suggesting the presence of strong corrections to scaling, likely still affecting the presented estimates of $1/\nu$ using Eq. (5e) in Table 1. Another exception are the $\gamma/\nu$ exponents from the fit for the peaks of the magnetic susceptibility with subtraction. It is well known that using the magnetic susceptibility with subtraction (especially of the finite-size type $\langle|m|\rangle$) is subject to strong corrections (see, e.g., Refs. [54, 55, 56, 57]). The disconnected form $\chi^{\prime}$ for $\beta\leq\beta_{c}$ evaluated at $\beta=\beta_{c}$ commonly shows better scaling, as is also the case here. Nonetheless, even the $\gamma/\nu$ values from $\chi_{L}^{\prime}(\beta_{c}$) are about two to three standard deviations away from $1.75$. For $J_{2}=-0.21$ and 0.22-0.22 we use our $\beta_{c}$ estimates of Ref. [24], i.e., $\beta_{c}=3.2572(5)$ and $4.2702(6)$. For $J_{2}=-0.23$ we use $6.346(3)$ from the scaling of the location of $\mathrm{d}\hskip-0.50003pt\ln|m|$, i.e., $\beta_{\max}^{(\mathrm{d}\hskip-0.40974pt\ln|m|)}$. Similarly, evaluating the magnetization at $\beta_{\max}^{(\mathrm{d}\hskip-0.40974pt\ln|m|)}$ and at $\beta_{c}$ gives different estimates for the exponent ratio $\beta/\nu$. Table 1 also lists the results from the partition function zeros studied in Ref. [24] for $J_{2}=-0.21$ and 0.22-0.22, which, within error bars, are compatible with the values for $\gamma/\nu$, $1/\nu$, and $\beta_{c}$ obtained here.